OpenStudy (anonymous):

rationalize the denominator of 5/2sqrt.of3

OpenStudy (anonymous):

\[\frac{5}{2\sqrt{3}}=\frac{5}{2\sqrt{3}}\times \frac{\sqrt{3}}{\sqrt{3}}\] \[\large =\frac{5\sqrt{3}}{2(3)}\] \[\large =\frac{5\sqrt{3}}{6}\]

OpenStudy (anonymous):

You have to just get rid of the square root int he denominator when rationalising. And you know when you don't want a square root sign, you multiply the square root of the number with a square root of the same number. That's all to it.

OpenStudy (anonymous):

in the*

OpenStudy (anonymous):

Thanks so much. :)